OpenStudy (anonymous):

Decide whether or not the ordered pair is a solution of the equation. 2x - 4y = 22 (3,4) A. Yes B. No

OpenStudy (anonymous):

I think yes if (3,4) was a solution 2x−4y=22 2×(3)−4×(4) would equal =22 ?

OpenStudy (ajprincess):

2×(3)−4×(4)=6-16=-10

OpenStudy (ajprincess):

Nw vat do u thnk is the answer?

OpenStudy (anonymous):

No solution

OpenStudy (ajprincess):

yup u r right:)